# Towards Non-Stationary Time Series Forecasting with Temporal Stabilization and Frequency Differencing

### v3

@@ Abstract
Time series forecasting is critical for decision-making across dynamic domains such as energy, finance, transportation, and cloud computing. However, real-world time series often exhibit non-stationarity, including temporal distribution shifts and spectral variability, which pose significant challenges for long-term time series forecasting. In this paper, we propose DTAF, a dual-branch framework that addresses non-stationarity in both the temporal and frequency domains. For the temporal domain, the Temporal Stabilizing Fusion (TFS) module employs a non-stationary mix of experts (MOE) filter to disentangle and suppress temporal non-stationary patterns while preserving long-term dependencies. For the frequency domain, the Frequency Wave Modeling (FWM) module applies frequency differencing to dynamically highlight components with significant spectral shifts. By fusing the complementary outputs of TFS and FWM, DTAF generates robust forecasts that adapt to both temporal and frequency domain non-stationarity. Extensive experiments on real-world benchmarks demonstrate that DTAF outperforms state-of-the-art baselines, yielding significant improvements in forecasting accuracy under non-stationary conditions. All codes are available at https://github.com/PandaJunk/DTAF.
@@ Introduction
Multivariate time series forecasting aims to predict future values of multiple channels based on their historical data (dl; cheng2023weakly; wang2023hierarchical). Traditional methods such as RNN (RNN; ts-RNN) and LSTM (lstm) capture time dependencies, but as the sequence length increases, the historical patterns will be lost. To address these challenges, deep learning approaches such as DeepAR (DeepAR) have emerged, utilizing autoregressive probabilistic models to enhance robustness and accuracy in forecasting. Recent advancements have introduced Graph Neural Networks (GNNs) (enhanceNet; gnn; tpgnn) to model spatial dependencies in time series. For example, in tasks such as energy forecasting (ELC), finance prediction (wang2025plug), and traffic prediction (Rela-traffic), GNNs are capable of capturing both temporal information and spatial correlations between variables, thereby improving forecasting accuracy. Simultaneously, Transformer-based architectures use attention mechanisms to model long-range dependencies, such as PatchTST (PatchTST), Pathformer (pathformer).
@@
The core of design is TFS and FWM, which are designed to model non-stationarity in the temporal and frequency domains. TFS utilizes a Non-stationary MoE Filter to dynamically extract and remove non-stationary patterns and then uses the Temporal Fusion module to model long-term dependency based on the stationary residual. Complementing the temporal modeling, FWM addresses frequency non-stationarity by applying differencing in the frequency domain, highlighting components with the most significant spectral changes. These enable DTAF to adapt to temporal and frequency dynamics simultaneously, leading to more accurate forecasting under non-stationary conditions. To maintain clarity and focus in the main text, we include essential definitions and task formulations in Appendix A.
@@
For notational clarity, we describe our method using a univariate time series $\mathbf{X}\in\mathbb{R}^{T}$, where $T$ denotes the input sequence length (extension to multivariate cases is straightforward). We first apply instance normalization to standardize the input statistics, mitigating distribution shift. Recognizing that individual time steps often lack meaningful context, we employ a patching strategy that segments the sequence into NN overlapping patches $\{\mathbf{X}^{1},\mathbf{X}^{2},\cdots,\mathbf{X}^{N}\}$, where each patch captures local temporal patterns. These patches are then projected into a high-dimensional latent space through patch embedding, yielding the embeddings $\{\mathbf{X}_{\text{patch}}^{1},\mathbf{X}_{\text{patch}}^{2},\cdots,\mathbf{X}_{\text{patch}}^{N}\}$ for subsequent processing, where $\mathbf{X}_{\text{patch}}^{i}\in\mathbb{R}^{d}$ represents $i$-th patch embedding, and $d$ denotes the embedding dimension.
@@
where $\mathbf{X}_{\mathrm{out}}\in\mathbb{R}^{F}$, $F$ is the prediction length of time series. This module adaptively selects the required features from both the temporal and frequency domains to produce the final prediction $\mathbf{X}_{\mathrm{out}}$.
@@
Table 2: Ablation Study. W/O TFS, W/O FWM, represent removing the TFS and FWM, and W/ Attention represents replacing Dual-branch Attention with Cross Attention.

### v5

@@ Abstract
Time series forecasting is critical for decision-making across dynamic domains such as energy, finance, transportation, and cloud computing. However, real-world time series often exhibit non-stationarity, including temporal distribution shifts and spectral variability, which pose significant challenges for long-term time series forecasting. In this paper, we propose DTAF, a dual-branch framework that addresses non-stationarity in both the temporal and frequency domains. For the temporal domain, the Temporal Stabilizing Fusion (TFS) module employs a non-stationary mix of experts (MOE) filter to disentangle and suppress temporal non-stationary patterns while preserving long-term dependencies. For the frequency domain, the Frequency Wave Modeling (FWM) module applies frequency differencing to dynamically highlight components with significant spectral shifts. By fusing the complementary outputs of TFS and FWM, DTAF generates robust forecasts that adapt to both temporal and frequency domain non-stationarity. Extensive experiments on real-world benchmarks demonstrate that DTAF outperforms state-of-the-art baselines, yielding significant improvements in forecasting accuracy under non-stationary conditions. All codes are available at https://github.com/decisionintelligence/DTAF.
@@ Introduction
Multivariate time series forecasting aims to predict future values of multiple channels based on their historical data (qiu2025comprehensive; cheng2023weakly; wang2023hierarchical). Traditional methods such as RNN (RNN; ts-RNN) and LSTM (lstm) capture time dependencies, but as the sequence length increases, the historical patterns will be lost. To address these challenges, deep learning approaches such as DeepAR (DeepAR) have emerged, utilizing autoregressive probabilistic models to enhance robustness and accuracy in forecasting. Recent advancements have introduced Graph Neural Networks (GNNs) (enhanceNet; gnn; tpgnn) to model spatial dependencies in time series. For example, in tasks such as energy forecasting (ELC; tian2024Air) , finance prediction (wang2025plug), and traffic prediction (Rela-traffic), GNNs are capable of capturing both temporal information and spatial correlations between variables, thereby improving forecasting accuracy. Simultaneously, Transformer-based architectures use attention mechanisms to model long-range dependencies, such as PatchTST (PatchTST), Pathformer (pathformer).
@@
The core of design is TFS and FWM, which are designed to model non-stationarity in the temporal and frequency domains. TFS utilizes a Non-stationary MoE Filter to dynamically extract and remove non-stationary patterns and then uses the Temporal Fusion module to model long-term dependency based on the stationary residual. Complementing the temporal modeling, FWM addresses frequency non-stationarity by applying differencing in the frequency domain, highlighting components with the most significant spectral changes. These enable DTAF to adapt to temporal and frequency dynamics simultaneously, leading to more accurate forecasting under non-stationary conditions. To maintain clarity and focus in the main text, we include essential definitions and task formulations in Appendix B.
@@
For notational clarity, we describe our method using a univariate time series $\mathbf{X}\in\mathbb{R}^{T}$, where $T$ denotes the input sequence length (extension to multivariate cases is straightforward). We first apply instance normalization to standardize the input statistics, mitigating distribution shift. Recognizing that individual time steps often lack meaningful context, we employ a patching strategy that segments the sequence into $N$ overlapping patches $\{\mathbf{X}^{1},\mathbf{X}^{2},\cdots,\mathbf{X}^{N}\}$, where each patch captures local temporal patterns. These patches are then projected into a high-dimensional latent space through patch embedding, yielding the embeddings $\{\mathbf{X}_{\text{patch}}^{1},\mathbf{X}_{\text{patch}}^{2},\cdots,\mathbf{X}_{\text{patch}}^{N}\}$ for subsequent processing, where $\mathbf{X}_{\text{patch}}^{i}\in\mathbb{R}^{d}$ represents $i$-th patch embedding, and $d$ denotes the embedding dimension.
@@
where $\mathbf{X}_{\mathrm{out}}\in\mathbb{R}^{F}$, $F$ is the prediction length of time series. This module adaptively selects the required features from the two domains to produce the final prediction $\mathbf{X}_{\mathrm{out}}$.
@@
Table 2: Ablation Study. W/O TFS, W/O FWM, represent removing the TFS and FWM, and W/ Attention represents replacing Dual-branch Attention with Cross Attention. Results averaged over 3 seeds; std $\leq$ ± 0.0058.
@@ Conclusion
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (62372179, 62406112).